OpenStudy (anonymous):

solve equation: 6(-x-5)= -4(x-3)-x

OpenStudy (anonymous):

i just need to check my answer now, i think i got it

OpenStudy (anonymous):

x=42

OpenStudy (anonymous):

i got x=32

OpenStudy (anonymous):

-6x-30=-4x+12-x -6x-30=-5x+12 -30=x+12 -42=x

OpenStudy (anonymous):

ahh i see where i messed up :l

OpenStudy (anonymous):

oh yeah, i lost minus

OpenStudy (anonymous):

can you show me the check to your solution? i messed that up too. *sigh*

OpenStudy (anonymous):

i know but i have to show my check

OpenStudy (anonymous):

6(-x-5)= -4(x-3)-x 6(42-5)=-4(-42-3)-42 6(38)=-4(-45)-42 Do you think you can do it from there?

OpenStudy (anonymous):

6(38)=-4(-45)-42 doesn't come out equal.. does it?

OpenStudy (anonymous):

@katylon this is what you do to check it: 6(42-5)=-4(-42-3)+42